# MA-DV2F: A Multi-Agent Navigation Framework using Dynamic Velocity Vector Field

### v1

@@ I INTRODUCTION
The task of multi-agent navigation has attracted widespread attention in recent years due to myriad applications in areas such as search and rescue missions [1], area exploration [2], pickup and delivery services [3], warehouses [4], self-driving cars [5] etc. The task of multi-agent path finding/navigation involves simultaneously directing a group of vehicles from their initial position to their desired destination while avoiding collisions with other agents. The task is known to be NP-hard even in the discrete setting [6]. An ideal algorithm must find the optimal solution in limited time. This leads to contradictory goals, since determining the optimal solution requires searching a larger solution space, which necessitates more time. In structured environments such as indoor spaces, prior knowledge and understating of the layout impose constraints, that can reduce the solution search space. In unstructured environments, there are no such constraints. This allows an algorithm the flexibility to find a solution. However, since the search space is much larger, there is no guarantee that the solution found is optimal. The problem is further exacerbated when the search space in continuous and agents are non-holomonic vehicles. The constraints arising from the vehicle kinematics add to the complexity.
@@
There have been various techniques and heuristics attempting to find (near-)optimal trajectories for multiple agents. The methods can be divided into two primary categories: 1) Learning based data driven methods [7, 8] and 2) Search/optimization based methods [9, 10]. Learning based algorithms involve training a neural network on data, with the understating that the network will generalize at inference time. The training data should encompass all the situations that the model is expected to encounter at test time. This necessitates a large amount of training samples. The greatest challenge with large training samples is determining the corresponding supervised labels for training; that might be too tedious to obtain. In contrast to supervised learning, an alternate would be to train using reinforcement learning (RL) where the model explores the environment and acquires rewards or penalties depending on the actions taken [11]. The model then exploits this experience to execute the correct control actions at test time. However, RL algorithms tend to be more sample inefficient than supervised methods. In contrast, optimization [12] or search based [13] methods involves simultaneously optimizing trajectories for multiple vehicles. As the number of vehicles are added, the complexity of the optimization/search becomes intractable making it infeasible for scaling to a large number of vehicles [14, 15].
@@
Figure 1: Shows the pipeline for MA-DV2F. The state of all vehicles at time $t$ is used to create the dynamic velocity vector field (DV2F) from which the reference control commands are determined. These commands are in turn used to determine the next state at time $t+1$ using the kinematic motion model. Note that the reference control commands can optionally be used to train a GNN in a self-supervised manner (indicated by dotted arrows). Meanwhile, the DV2F is shown for the black ego-vehicle with 2 other vehicles (red & blue) and 2 obstacles (green & brown) in the scene. The target state for the black ego-vehicle is shown by the dotted rectangular with a solid arrow at the top of the map. The dotted circle around the target state is the parking region. The black arrows indicate the ideal reference orientation as unit vectors which the ego-vehicle should attain at each different position on the map. The colored dotted circles around other vehicles and obstacles show the collision avoiding regions for the black ego-vehicle. Each black arrow in these regions is composed of an attractive target approaching component (gray arrow) and a repulsive collision avoidance component (colored arrow). Note that due to kinematic constraints, the ideal orientation might not be attainable. The shaded wedge in front of the ego-vehicle shows the reachable angle closest to the ideal. An example of the dynamic velocity vector field is shown in the project page: https://yininghase.github.io/MA-DV2F/#VD. Note that DV2F for the neighboring red & blue vehicles is likewise created separately.
@@
In this paper, we propose Multi-Agent Dynamic Velocity Vector Field (MA-DV2F), which generates vectors for the reference speed and orientation vectors for each position on the map for every vehicle. The vehicles then just need to follow in the direction of their respective velocity vector fields to successfully reach their destinations. The vector field for each vehicle is generated independently and can be adapted dynamically depending on the vehicle’s proximity to other agents (neighbouring vehicles or obstacles). Decoupling reduces the complexity of adding vehicles & allows for parallel generation of DV2F of each vehicle, thereby increasing the throughput. An added benefit of our approach is that the generated DV2F can be used to train a learning based graphical neural network (GNN) in a self-supervised manner. This self-supervised learning based approach neither requires tedious labeling of data nor necessitates sample inefficient environment exploration. We test our framework under challenging collision prone environments. A scenario is regarded to be challenging if trajectories of different vehicles considered independently from other agents intersect at multiple places at the same time. This would necessitate a collision avoidance maneuver for safe navigation towards the target.
@@
Figure 1 shows the pipeline for both the MA-DV2F (left branch, solid arrows) and optional training of the the self-supervised GNN counterpart (right branch, dotted arrows). The input is a continuous state representation of all the multiple vehicles and the outputs are the corresponding continuous control variables. The vehicles are non-holonomic with rotation radius determined by the kinematic model.
@@ II Related Work
[16] proposed using Artificial Potential Fields (APF) for trajectory planning. An agent is subjected to an attractive potential force towards the target which serves as the sink and a repulsive potential force away from obstacles [17, 18]. However, a common problem with such methods is their propensity to get stuck in local minima [19, 20] when the attractive force from the target is cancelled out by the repulsive force arising from an another agent for e.g. when the ego-vehicle is symmetrically aligned with other agents to the target [21] and thus leading to a bottleneck situation. We break such bottlenecks by enforcing the vehicles to move in the counterclockwise direction.
@@
CL-MAPF [9] uses a Body Conflict Tree to describe agent collision scenarios as spatiotemporal constraints. It then applies a Hybrid-State A* Search algorithm to generate paths satisfying both kinematic and spatiotemporal constraints of the vehicles. However, under challenging test scenarios with vehicles crowding together, the algorithm takes long to search for a solution and can easily time out. To find feasible solution for large-scale multi-vehicle trajectory planning, CSDO [10] first searches for a coarse initial guess using a large search step. Then, the Decentralized Quadratic Programming is implemented to refine this guess for minor collisions. GCBF+ [8] based on GCBF [7] aims to provide safety guarantees utilizing control barrier functions (CBF). A graphical Neural Network is trained in an RL styled manner to learn agent control policy.
@@ III-A Problem Description
We aim to solve the task of multi-agent navigation in unconstrained environments. Given $N_{veh}$ dynamic vehicles and $N_{obs}$ static obstacles in the scene, the task is to ensure each vehicle reaches its desired destination while avoiding collision with other agents. The state vector for vehicle $i$ ($1\leq i\leq N_{veh}$) at current time $t$ can be represented as $\mathbf{s}_{t}^{({i}_{veh})}=[x_{t},y_{t},\theta_{t},v_{t},x_{tar},y_{tar},\theta_{tar}]^{T}$, where $x_{t}$ and $y_{t}$ shows the position, $\theta_{t}$, the orientation and $v_{t}$ the speed at current time $t$. Meanwhile, $x_{tar}$ and $y_{tar}$ are co-ordinates of the target position, and $\theta_{tar}$ is the target orientation. Each vehicle is controlled by a control vector $\mathbf{c}_{t}^{({i}_{veh})}=[p_{t},\varphi_{t}]^{T}$, where $p_{t}\in[-P,P]$ and $\varphi_{t}\in[-\Phi,\Phi]$ are the pedal acceleration and steering angle, limited in magnitude by $P$ and $\Phi$ respectively. The obstacle $i$ ($1\leq i\leq N_{obs}$) is represented by a state vector $\mathbf{s}^{({i}_{obs})}=[x,y,r]^{T}$, where $x$ and $y$ is the position, and $r$ is the radius of the circle circumscribing all vertices of the obstacle. The kinematics of the vehicles is modeled using the bicycle model [28].
@@
It describes how the equations of motion can be updated in time increments of $\Delta t$ assuming no slip condition, valid under low or moderate vehicle speed $v$ when making turns. Note that $\beta$ and $\gamma$ are the tuneable hyper-parameters.
@@
  • Estimation of reference orientation of ego-vehicle for every position in the map
  • Estimation of reference speed for the corresponding positions.
The orientation which a vehicle should possess at any position on the map is referred to as the Reference Orientation and is represented as a unit vector. The black arrows in Figure 1 show the reference orientation for the black ego vehicle at different positions in the map. It can be seen that the arrows attract the ego-vehicle towards its target while repelling it away from the other vehicles and obstacles. Note that the current orientation of the ego-vehicle is not aligned with the reference orientation at its existing position. Therefore, we ought to find the control variables which align the ego-vehicle with the reference orientation. Apart from the reference orientation at each position in the map, the ego-vehicle also has a reference speed which should be attained by the control variables. Lastly, note that a separate reference orientation map also exists for the blue and red vehicles shown in Figure 1.
@@ III-B Estimation of Reference Orientation
We first define some frequently used functions and vectors before we begin discussion of reference orientation estimation.
The vector function $\mathbf{f_{uni}}(\mathbf{a})$ is the function which takes a non-zero vector $\mathbf{a}$ as input and divides by its magnitude to convert it into a unit vector.
@@
  • the position $(x^{(j)}_{obs},y^{(j)}_{obs})$ of the static obstacle $j$ as $\mathbf{X}^{(i)}_{obs_{j}}=[x^{(j)}_{obs}-x^{(i)}_{t+1},y^{(j)}_{obs}-y^{(i)}_{t+1}]^{T}$
@@
where $r_{p}$ is the parking threshold (black dotted circle in Figure 1) and $0.5\cdot v^{2}_{d}$ is the marginal parking threshold (shaded blue region in Figure 1). $v_{d}$ is the default reference speed, estimation of which is explained in Subsection III-C. $\epsilon_{p}$ is the threshold above which an additional term is introduced when the vehicle is within the parking threshold.
Equation 2 shows that when the ego-vehicle $i$ is far away from the parking threshold, the reference orientation is in the direction of $\mathbf{X}^{(i)}_{tar}$. However, as the ego-vehicle approaches the target and the velocity is high, then the vehicle might overshoot the target and enter the shaded marginal parking threshold region. In this case the direction of the orientation is flipped using $f_{sgn}({\mathbf{X}_{tar}^{(i)}}^{T}.\mathbf{U}^{(i)}_{t})$. This is to prevent the vehicle from moving in circles. A detailed explanation of this is given in the supplementary in the project page.
@@
Besides reaching the target, the ego vehicle $i$ should also avoid collision on its way to the target. This is realized by the collision avoiding component $\mathbf{u}^{(i)}_{coll}$ which comprises of collision avoidance between the ego vehicle $i$ and either static obstacle $j$ ($\mathbf{u}^{(i)}_{obs_{j}}$) or with another vehicle $j$ ($\mathbf{u}^{(i)}_{veh_{j}}$).
The equation for determining ($\mathbf{u}^{(i)}_{obs_{j}}$) is given by:
@@
$\begin{split}\mathbf{u}^{(i)}_{obs_{j}}&=\begin{cases}\mathbf{f_{uni}}(\mathbf{X}^{(i)}_{obs_{j}})\cdot\alpha^{(i)}_{obs_{j}}+\mathbf{R}^{(i)}_{obs_{j}}\cdot\beta^{(i)}_{obs_{j}}&\alpha^{(i)}_{obs_{j}}\leq 0\\ \mathbf{0}&\text{otherwise}\\ \end{cases}\\ \alpha^{(i)}_{obs_{j}}&=\|\mathbf{X}^{(i)}_{obs_{j}}\|_{2}-r_{obs}^{(j)}-r_{veh}-(r_{c}+|v^{(i)}_{t}|)\\ \beta^{(i)}_{obs_{j}}&=f_{pos}({\mathbf{X}^{(i)}_{tar}}^{T}\cdot\mathbf{X}_{obs_{j}}^{(i)})\cdot(\|\mathbf{X}^{(i)}_{obs_{j}}\|_{2}-r_{obs}^{(j)})\\ \mathbf{R}^{(i)}_{obs_{j}}&=\mathbf{f_{uni}}(\mathbf{Z}\times\mathbf{X}^{(i)}_{obs_{j}})\\ \end{split}$ (3)
@@
where $r^{(j)}_{obs}$ is the radius of the static object $j$, $r_{veh}$ is the radius of the smallest circle enclosing the ego-vehicles, $r_{c}$ is the static part, while $|v^{(i)}_{t}|$ is the speed-based dynamic margin for collision avoidance between ego vehicle $i$ and obstacle $j$. Higher the vehicle speed $|v^{(i)}_{t}|$, larger is this margin $r_{c}+|v^{(i)}_{t}|$. When the distance between the ego vehicle $i$ and static object $j$ ($\|\mathbf{X}^{(i)}_{obs_{j}}\|_{2}-r_{obs}^{(j)}-r_{veh}$) is below the collision avoidance margin $r_{c}+|v^{(i)}_{t}|$, then $\alpha^{(i)}_{obs_{j}}\leq 0$, and the reference orientation is modified to prevent collision with the static obstacle. Under this condition, the first term $\mathbf{f_{uni}}(\mathbf{X}^{(i)}_{obs_{j}})\cdot\alpha^{(i)}_{obs_{j}}$ will guide the vehicle to drive away from the static obstacle. However, driving away is not enough as this might cause a bottleneck in cases where the obstacle is symmetrically collinear between the ego-vehicle and its target. We would additionally like the ego-vehicle to drive around the obstacle to reach the target for which the term $\mathbf{R}^{(i)}_{obs_{j}}\cdot\beta^{(i)}_{obs_{j}}$ is relevant. If the ego-vehicle is between the obstacle and target then $\beta^{(i)}_{obs_{j}}=0$ (since $f_{pos}({\mathbf{X}^{(i)}_{tar}}^{T}\cdot\mathbf{X}_{obs_{j}}^{(i)})=0$) and hence there is no need for the vehicle to drive around the obstacle. However, if that is not the case, then an additional component is added whose direction is given by $\mathbf{R}^{(i)}_{obs_{j}}$ and magnitude ($\beta^{(i)}_{obs_{j}}$) is proportional to how far the vehicle is away from the obstacle. $\mathbf{R}^{(i)}_{obs_{j}}$ is given as the cross product between $\mathbf{Z}$ and $\mathbf{X}^{(i)}_{obs_{j}}$ with a zero being appended to the third dimension of $\mathbf{X}^{(i)}_{obs_{j}}$ which originally lies in the 2D space. The vector resulting from the cross-product is perpendicular to $\mathbf{X}_{obs_{j}}^{(i)}$ which causes the vehicles to move in the counterclockwise direction around the obstacle as shown in Figure 2.
@@
Figure 2: Shows the constituents of the reference orientation vector $\mathbf{u}_{ref}$ near an obstacle. It comprises of a target approaching ($\mathbf{u}_{tar}$) and the collision avoiding ($\mathbf{u}_{obj}$) components. $\mathbf{u}_{obj}$ includes $\mathbf{v}_{n}=\mathbf{f_{uni}}(\mathbf{X}^{(i)}_{obs_{j}})\cdot\alpha^{(i)}_{obs_{j}}$ guiding the vehicle to drive away from the obstacle and $\mathbf{v}_{t}=\mathbf{R}^{(i)}_{obs_{j}}\cdot\beta^{(i)}_{obs_{j}}$ leading the vehicle to go around the obstacle. When the ego-vehicle is between the obstacle and target, it is not necessary for it to go around the obstacle and thus $\mathbf{v}_{t}=\mathbf{0}$. The formulation to calculate $\mathbf{u}_{obj}$ is described in Equation 3.
@@
Likewise, the component for avoiding collision between the ego vehicle $i$ and another vehicle $j$ ($\mathbf{u}^{(i)}_{veh_{j}}$) is similar to Equation 3 except that the static obstacle radius $r_{obs}^{(j)}$ will be replaced by the other vehicle’s radius $r_{veh}$ in the $\alpha^{(i)}_{veh_{j}}$ term. Secondly, the speed based dynamic margin is $|v^{(j)}_{t}|+|v^{(i)}_{t}|$ rather than just $|v^{(i)}_{t}|$.
@@
$\mathbf{u}^{(i)}_{coll}=\sum_{j=1}^{N_{obs}}\mathbf{u}^{(i)}_{obs_{j}}+\sum_{j=1,j\neq i}^{N_{veh}}\mathbf{u}^{(i)}_{veh_{j}}$ (4)
Finally, the ideal reference orientation vector $\mathbf{\hat{u}}^{(i)}_{t+1}$ is:
@@
From this, the corresponding ideal reference orientation angle $\hat{\theta}^{(i)}_{t+1}$ for ego vehicle $i$ can be calculated by applying $arctan2$ to $\mathbf{\hat{u}}^{(i)}_{t+1}$. However, kinematic constraints arising from the motion model Equation 1 may prevent the vehicle from immediately attaining the ideal reference orientation $\hat{\theta}^{(i)}_{t+1}$ in the next time step. Therefore, we instead use $\theta^{(i)}_{t+1}$ referred to as the real orientation. It is the reachable orientation closest to $\hat{\theta}^{(i)}_{t+1}$. The unit vector corresponding to this real reference orientation angle $\theta^{(i)}_{t+1}$ for ego vehicle $i$ is $\mathbf{u}^{(i)}_{t+1}=[\cos(\theta^{(i)}_{t+1}),\sin(\theta^{(i)}_{t+1})]^{T}$.

### v2

@@ I INTRODUCTION
The task of multi-agent navigation has attracted widespread attention in recent years due to myriad applications in areas such as search and rescue missions [1], area exploration [2], pickup and delivery services [3], warehouses [4], self-driving cars [5] etc. The task of multi-agent path finding/navigation involves simultaneously directing a group of vehicles from their initial position to their desired destination while avoiding collisions with other agents. The task is known to be NP-hard even in the discrete setting [6]. An ideal algorithm must find the optimal solution in limited time. This leads to contradictory goals, since determining the optimal solution requires searching a larger solution space, which necessitates more time. In structured environments such as indoor spaces, prior knowledge and understating of the layout impose constraints, that can reduce the solution search space. In unstructured environments, there are no such constraints. This allows an algorithm the flexibility to find a solution. However, since the search space is much larger, there is no guarantee that the solution found is optimal. The problem is further exacerbated when the search space is continuous and agents are non-holomonic vehicles. The constraints arising from the vehicle kinematics add to the complexity.
@@
There have been various techniques and heuristics attempting to find (near-)optimal trajectories for multiple agents. The methods can be divided into two primary categories: 1) Learning based data driven methods [7, 8] and 2) Search/optimization based methods [9, 10]. Learning based algorithms involve training a neural network on data, with the understanding that the network will generalize at inference time. The training data should encompass all the situations that the model is expected to encounter at test time. This necessitates a large amount of training samples. The greatest challenge with large training samples is determining the corresponding supervised labels for training; that might be too tedious to obtain. In contrast to supervised learning, an alternate would be to train using reinforcement learning (RL) [11], where the model explores the environment and acquires rewards or penalties depending on the actions taken. The model then exploits this experience to execute the correct control actions at test time. However, RL algorithms tend to be more sample inefficient than supervised methods. In contrast, optimization [12] or search based [13] methods involves simultaneously optimizing trajectories for multiple vehicles. As the number of vehicles are added, the complexity of the optimization/search becomes intractable making it infeasible for scaling to a large number of vehicles [14, 15].
@@
Figure 1: Shows the pipeline for MA-DV2F. The state of all vehicles at time $t$ is used to create the dynamic velocity vector field (DV2F) from which the reference control commands are determined. These commands are in turn used to determine the next state at time $t+1$ using the kinematic motion model. Note that the reference control commands can optionally be used to train a GNN in a self-supervised manner (indicated by dotted arrows). Meanwhile, the DV2F is shown for the black ego-vehicle with 2 other vehicles (blue & maroon) and 2 obstacles (green & brown) in the scene. The target state for the black ego-vehicle is shown by the dotted rectangle with a solid arrow at the top of the map. The dotted circle around the target state is the parking region. The black arrows indicate the ideal reference orientation as unit vectors which the ego-vehicle should attain at each different position on the map. The colored dotted circles around other vehicles and obstacles show the collision avoiding regions for the black ego-vehicle. Each black arrow in these regions is composed of an attractive target approaching component (gray arrow) and a repulsive collision avoidance component (colored arrow). Note that due to kinematic constraints, the reference orientation ($\mathbf{u}_{\mathbf{ref}}$) might not be attainable for the ego-vehicle at its existing location. The shaded wedge in front of the ego-vehicle shows the region of reachable orientation at the next time step. The real orientation ($\mathbf{u}_{\mathbf{real}}$) is therefore the physically attainable orientation by the ego-vehicle that is closest to the reference.
An example of the dynamic velocity vector field is shown in the project page: https://yininghase.github.io/MA-DV2F/#VD. Note that DV2F for the neighboring blue & maroon vehicles is likewise created separately (not shown here).
@@
In this paper, we propose Multi-Agent Dynamic Velocity Vector Field (MA-DV2F), which generates vectors for the orientation and reference speed for every vehicle on the map. The vehicles then just need to follow in the direction of their respective velocity vector fields to successfully reach their destinations. The vector field for each vehicle is generated independently and can be adapted dynamically depending on the vehicle’s proximity to other agents (neighbouring vehicles or obstacles). Decoupling reduces the complexity of adding vehicles & allows for parallel generation of DV2F of each vehicle, thereby increasing the throughput. An added benefit of our approach is that the generated DV2F can be used to train a learning based graphical neural network (GNN) in a self-supervised manner. This self-supervised learning based approach neither requires tedious labeling of data nor necessitates sample inefficient environment exploration. We test our framework under challenging collision prone environments. A scenario is regarded to be challenging if trajectories of different vehicles considered independently from other agents intersect at multiple places at the same time. This would necessitate a collision avoidance maneuver for safe navigation towards the target.
@@
Fig. 1 shows the pipeline for both the MA-DV2F (left branch, solid arrows) and optional training of the the self-supervised GNN counterpart (right branch, dotted arrows). The input is a continuous state representation of all the multiple vehicles and the outputs are the corresponding continuous control variables. The vehicles are non-holonomic with rotation radius determined by the kinematic model.
@@ II Related Work
[16] proposed using Artificial Potential Fields (APF) for trajectory planning. An agent is subjected to an attractive potential force towards the target which serves as the sink and a repulsive potential force away from obstacles [17, 18]. However, a common problem with such methods is their propensity to get stuck in local minima [19, 20] when the attractive force from the target is cancelled out by the repulsive force arising from an another agent for e.g. when the ego-vehicle is symmetrically aligned with other agents to the target [21] and thus leading to a bottleneck situation. We break such bottlenecks by enforcing the vehicles to move in the clockwise direction.
@@
CL-MAPF [9] uses a Body Conflict Tree to describe agent collision scenarios as spatiotemporal constraints. It then applies a Hybrid-State A* Search algorithm to generate paths satisfying both kinematic and spatiotemporal constraints of the vehicles. However, under challenging test scenarios with vehicles crowding together, the algorithm takes long to search for a solution and can easily time out. To find feasible solution for large-scale multi-vehicle trajectory planning, CSDO [10] first searches for a coarse initial guess using a large search step. Then, the Decentralized Quadratic Programming is implemented to refine this guess for minor collisions. GCBF+ [8] based on GCBF [7] aims to provide safety guarantees utilizing control barrier functions (CBF). A Graphical Neural Network is trained to learn agent control policy.
@@ III-A Problem Description
We aim to solve the task of multi-agent navigation in unconstrained environments. Given $N_{veh}$ dynamic vehicles and $N_{obs}$ static obstacles in the scene, the task is to ensure each vehicle reaches its desired destination while avoiding collision with other agents. The state vector for ego-vehicle $i$ ($1\leq i\leq N_{veh}$) at current time $t$ can be represented as $\mathbf{s}_{t}^{({i})}=[x_{t},y_{t},\theta_{t},v_{t},x_{tar},y_{tar},\theta_{tar}]^{T}$, where $x_{t}$ and $y_{t}$ shows the position, $\theta_{t}$, the orientation and $v_{t}$ the speed at current time $t$. Meanwhile, $x_{tar}$ and $y_{tar}$ are co-ordinates of the target position, and $\theta_{tar}$ is the target orientation. Each ego-vehicle is controlled by a control vector $\mathbf{c}_{t}^{({i}_{veh})}=[p_{t},\varphi_{t}]^{T}$, where $p_{t}\in[-P,P]$ and $\varphi_{t}\in[-\Phi,\Phi]$ are the pedal acceleration and steering angle, limited in magnitude by $P$ and $\Phi$ respectively. The obstacle $k$ ($1\leq k\leq N_{obs}$) is represented by a state vector $\mathbf{s}^{k}=[x,y,r]^{T}$, where $x$ and $y$ is the position, and $r$ is the radius of the circle circumscribing all vertices of the obstacle. The kinematics of the vehicles is modeled using the bicycle model [28].
@@
It describes how the equations of motion can be updated in time increments of $\Delta t$ assuming no slip condition, valid under low or moderate vehicle speed $v$ when making turns. Note that $\beta$ and $\gamma$ are the tuneable hyper-parameters modeling the environment friction and vehicle wheelbase respectively.
@@
  • Estimation of reference orientation of ego-vehicle for every position in the map (See Subsection III-B).
  • Estimation of reference speed for the corresponding positions. (See Subsection III-C).
The orientation which a vehicle should possess at any position on the map is referred to as the Reference Orientation and is represented as a unit vector. The black arrows in Fig. 1 show the reference orientation for the black ego vehicle at different positions in the map. It can be seen that the arrows attract the ego-vehicle towards its target while repelling it away from the other vehicles and obstacles. Note that the current orientation of the ego-vehicle is not aligned with the reference orientation at its existing position. Therefore, we ought to find the control variables which align the ego-vehicle with the reference orientation. Apart from the reference orientation at each position in the map, the ego-vehicle also has a reference speed which should be attained by the control variables. Lastly, note that apart from the black vehicle, a separate reference orientation map is also created for the blue and red vehicles present in Fig. 1 (not shown in the Figure).
@@ III-B Estimation of Reference Orientation
We first define some frequently used functions and vectors before we begin discussion of reference orientation estimation. The vector function $\mathbf{f_{uni}}(\mathbf{a})$ is the function which takes a non-zero vector $\mathbf{a}$ as input and divides by its magnitude to convert it into a unit vector.
@@
  • the position $(x^{(k)}_{obs},y^{(k)}_{obs})$ of the static obstacle $k$ as $\mathbf{X}^{(i)}_{obs_{k}}=[x^{(k)}_{obs}-x^{(i)}_{t+1},y^{(k)}_{obs}-y^{(i)}_{t+1}]^{T}$
@@
where $r_{p}$ is the parking threshold (black dotted circle in Fig. 1) and $0.5\cdot v^{2}_{d}$ is the marginal parking threshold (shaded blue region in Fig. 1). $v_{d}$ is the default reference speed, estimation of which is explained in Subsection III-C. $\epsilon_{p}$ is the threshold above which an additional term is introduced when the vehicle is within the parking threshold.
Equation 2 shows that when the ego-vehicle $i$ is far away from the parking threshold, the reference orientation is in the direction of $\mathbf{X}^{(i)}_{tar}$. However, as the ego-vehicle approaches the target and the velocity is high, then the vehicle might overshoot the target and enter the shaded marginal parking threshold region. In this case the direction of the orientation is flipped using $f_{sgn}({\mathbf{X}_{tar}^{(i)}}^{T}.\mathbf{U}^{(i)}_{t})$. This is to prevent the vehicle from moving in circles. A detailed explanation of this is given in the supplementary file on the project page: https://yininghase.github.io/MA-DV2F/supplementary.pdf.
@@
Besides reaching the target, the ego vehicle $i$ should also avoid collision on its way to the target. This is realized by the collision avoiding component $\mathbf{u}^{(i)}_{coll}$ which comprises of collision avoidance between the ego vehicle $i$ and either static obstacle $k$ ($\mathbf{u}^{(i)}_{obs_{k}}$) or with another vehicle $j$ ($\mathbf{u}^{(i)}_{veh_{j}}$).
The equation for determining ($\mathbf{u}^{(i)}_{obs_{k}}$) is given by:
@@
$\begin{split}\mathbf{u}^{(i)}_{obs_{k}}&=\begin{cases}\mathbf{f_{uni}}(\mathbf{X}^{(i)}_{obs_{k}})\cdot\alpha^{(i)}_{obs_{k}}+\mathbf{R}^{(i)}_{obs_{k}}\cdot\beta^{(i)}_{obs_{k}}&\alpha^{(i)}_{obs_{k}}\leq 0\\ \mathbf{0}&\text{otherwise}\\ \end{cases}\\ \alpha^{(i)}_{obs_{k}}&=\|\mathbf{X}^{(i)}_{obs_{k}}\|_{2}-r_{obs}^{(k)}-r_{veh}-(r_{c}+|v^{(i)}_{t}|)\\ \beta^{(i)}_{obs_{k}}&=f_{pos}({\mathbf{X}^{(i)}_{tar}}^{T}\cdot\mathbf{X}_{obs_{k}}^{(i)})\cdot(\|\mathbf{X}^{(i)}_{obs_{k}}\|_{2}-r_{obs}^{(k)})\\ \mathbf{R}^{(i)}_{obs_{k}}&=\mathbf{f_{uni}}(\mathbf{Z}\times\mathbf{X}^{(i)}_{obs_{k}})\\ \end{split}$ (3)
@@
where $r^{(k)}_{obs}$ is the radius of the static object $k$, $r_{veh}$ is the radius of the smallest circle enclosing the ego-vehicles, $r_{c}$ is the static component, while $|v^{(i)}_{t}|$ is the speed-based dynamic safety margin for collision avoidance between ego vehicle $i$ and obstacle $k$. Higher the vehicle speed $|v^{(i)}_{t}|$, larger is this margin $r_{c}+|v^{(i)}_{t}|$. When the distance between the ego vehicle $i$ and static object $k$ ($\|\mathbf{X}^{(i)}_{obs_{k}}\|_{2}-r_{obs}^{(k)}-r_{veh}$) is below the collision avoidance margin $r_{c}+|v^{(i)}_{t}|$, then $\alpha^{(i)}_{obs_{k}}\leq 0$, and the reference orientation is modified to prevent collision with the static obstacle. Under this condition, the first term $\mathbf{f_{uni}}(\mathbf{X}^{(i)}_{obs_{k}})\cdot\alpha^{(i)}_{obs_{k}}$ will guide the vehicle to drive away from the static obstacle. However, driving away is not enough as this might cause a bottleneck in cases where the obstacle is symmetrically collinear between the ego-vehicle and its target. We would additionally like the ego-vehicle to drive around the obstacle to reach the target for which the term $\mathbf{R}^{(i)}_{obs_{k}}\cdot\beta^{(i)}_{obs_{k}}$ is relevant. If the ego-vehicle is between the obstacle and target then $\beta^{(i)}_{obs_{k}}=0$ (since $f_{pos}({\mathbf{X}^{(i)}_{tar}}^{T}\cdot\mathbf{X}_{obs_{k}}^{(i)})=0$) and hence there is no need for the vehicle to drive around the obstacle. However, if that is not the case, then an additional component is added whose direction is given by $\mathbf{R}^{(i)}_{obs_{k}}$ and magnitude ($\beta^{(i)}_{obs_{k}}$) is proportional to how far the vehicle is away from the obstacle. $\mathbf{R}^{(i)}_{obs_{k}}$ is given as the cross product between $\mathbf{Z}$ and $\mathbf{X}^{(i)}_{obs_{k}}$ with a zero being appended to the third dimension of $\mathbf{X}^{(i)}_{obs_{k}}$ which originally lies in the 2D space. The vector resulting from the cross-product is perpendicular to $\mathbf{X}_{obs_{k}}^{(i)}$ which causes the vehicles to move in the clockwise direction around the obstacle as shown in Fig. 2.
@@
Figure 2: Shows the constituents of the reference orientation vector $\mathbf{u}_{ref}$ near an obstacle. It comprises of a target approaching ($\mathbf{u}_{tar}$) and the collision avoiding ($\mathbf{u}_{obs_{k}}$) components. $\mathbf{u}_{obs_{k}}$ includes $\mathbf{v}_{n}=\mathbf{f_{uni}}(\mathbf{X}^{(i)}_{obs_{k}})\cdot\alpha^{(i)}_{obs_{k}}$ guiding the vehicle to drive away from the obstacle and $\mathbf{v}_{t}=\mathbf{R}^{(i)}_{obs_{k}}\cdot\beta^{(i)}_{obs_{k}}$ leading the vehicle to go around the obstacle. When the ego-vehicle is between the obstacle and target, it is not necessary for it to go around the obstacle and thus $\mathbf{v}_{t}=\mathbf{0}$. The formulation to calculate $\mathbf{u}_{obs_{k}}$ is described in Equation 3.
@@
Likewise, the component for avoiding collision between the ego vehicle $i$ and another vehicle $j$ ($\mathbf{u}^{(i)}_{veh_{j}}$) is similar to Equation 3 except that the static obstacle radius $r_{obs}^{(k)}$ will be replaced by the other vehicle’s radius $r_{veh}$ in the $\alpha^{(i)}_{veh_{j}}$ term. Secondly, the speed based dynamic margin is $|v^{(j)}_{t}|+|v^{(i)}_{t}|$ rather than just $|v^{(i)}_{t}|$.
@@
$\mathbf{u}^{(i)}_{coll}=\sum_{k=1}^{N_{obs}}\mathbf{u}^{(i)}_{obs_{k}}+\sum_{j=1,j\neq i}^{N_{veh}}\mathbf{u}^{(i)}_{veh_{j}}$ (4)
Finally, the ideal reference orientation vector $\mathbf{\hat{u}}^{(i)}_{t+1}$ ($\mathbf{u_{ref}}$ in Fig. 1) is:
@@
From this, the corresponding ideal reference orientation angle $\hat{\theta}^{(i)}_{t+1}$ for ego vehicle $i$ can be calculated by applying $arctan2$ to $\mathbf{\hat{u}}^{(i)}_{t+1}$. However, kinematic constraints arising from the motion model (Equation 1) may prevent the vehicle from immediately attaining the ideal reference orientation $\hat{\theta}^{(i)}_{t+1}$ in the next time step. Therefore, we instead use $\theta^{(i)}_{t+1}$ referred to as the real orientation. It is the reachable orientation closest to $\hat{\theta}^{(i)}_{t+1}$. The unit vector corresponding to this real reference orientation angle $\theta^{(i)}_{t+1}$ for ego vehicle $i$ is $\mathbf{u}^{(i)}_{t+1}$ ($\mathbf{u_{real}}$ in Fig. 1) = $[\cos(\theta^{(i)}_{t+1}),\sin(\theta^{(i)}_{t+1})]^{T}$.